OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 704t. After how many seconds does the projectile take to reach its maximum height?

OpenStudy (anonymous):

please please help!!

OpenStudy (anonymous):

22 seconds

OpenStudy (anonymous):

howd you find that? i have to show work.. could u explain to me so i understand please for when i take the test

OpenStudy (anonymous):

take the derivative and set it equal to 0

OpenStudy (anonymous):

wait ? huh im very bad in math lol

OpenStudy (anonymous):

oh i guess you havent taken calculus then? well in that case you can find the vertex of the equation using the formula -b/2a

OpenStudy (anonymous):

no.. this is for my online class for pre-call

OpenStudy (anonymous):

just use the equation -b/2a to find the vertex so it will be your b= 704 and a = -16 so if you plug it in you will get the time to get max height

OpenStudy (anonymous):

so i got 22 as the max height, where do i go from there ..what do i do now?

OpenStudy (anonymous):

no the 22 is the amount of time needed to get to max height, 22 is your answer your done

OpenStudy (anonymous):

ohh, so how would i put this into work to get the full credit? im sorry lol im so stupid when it comes to math

OpenStudy (anonymous):

its fine everyone has their own strengths, so what i would do is write down the vertext formula which is -b/2a then i would plug in those numbers and get 22 i would also draw a rough sketch of the parabola also and show that the highest point is at 22 if you can that is i guess you can also explain that the vertext tells you the point where the parabola goes the other direction, so basically it is the highest or lowest point on the graph depending on the equation so in your case it is the highest point

OpenStudy (anonymous):

so bc the max height is 22 is takes 22 seconds to reach it?

OpenStudy (anonymous):

the 22 is the amount of time it takes to reach the max height, the max height is the number when you plug in 22 into your equation

OpenStudy (anonymous):

ohok

OpenStudy (anonymous):

hope you understand everything i said, and gl with your test

OpenStudy (anonymous):

yeah kinda, for the most part. thank you!

OpenStudy (anonymous):

haha alright np glad to help

OpenStudy (anonymous):

this is what i got? is this wrong then t = -b/(2a) t = -740/(2(-16)) t = -740/(-32) t = 23.125 seconds

OpenStudy (anonymous):

its a typo, you wrote 704 when you asked the question so its should be like this: t= -b/2a t=-704/2(-16) t= 22

OpenStudy (anonymous):

wait.. what?

OpenStudy (anonymous):

you just made a type your wrote 740 instead of 704

OpenStudy (anonymous):

ohh opps

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

haha np